OpenStudy (anonymous):

A particle is moving with constant angular acceleration of 4 rad/s^2 in a circular path .At time T=0,PARTICLE was at rest . find the time at which the magnitude of centipetal acceleration and tangential acceleration are equal.

OpenStudy (anonymous):

@Herp_Derp @henpen @hei @shadowfiend

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

plz hlp me out

OpenStudy (anonymous):

@Herp_Derp hlp me out

OpenStudy (anonymous):

@Ruchi. I'm sorry I would've helped you out but honestly this is not my field.

OpenStudy (anonymous):

i hav cprrect it

OpenStudy (anonymous):

bt y c=0

OpenStudy (anonymous):

Actually, I'm probably wrong-let me corret myself

OpenStudy (anonymous):

As the particle starts at rest (at t=0, v=0, so only c=0 satisfies this)

OpenStudy (anonymous):

it hav correct it and get t=1/2 now tell me y c=0,what is c here reffer to?

OpenStudy (anonymous):

in think in case of differ. u hav refferd c.

OpenStudy (anonymous):

\(\alpha\) is constant:\[a_c=\frac{v^2}{r}\]\[v=r\omega=r\alpha t\]\[a_c=\frac{(r\alpha t)^2}{r}=r\alpha^2t^2\]\[a_t=r\alpha\]\[a_t=a_c \Leftrightarrow r\alpha=r\alpha^2t^2\]\[\frac{1}{\alpha}=t^2\]\[t=\sqrt{\frac{1}{\alpha}}\]

OpenStudy (anonymous):

henpen r u writng some thing or nt

OpenStudy (anonymous):

Do you understand my work @Ruchi ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Which part is confusing?

OpenStudy (anonymous):

\[v=r \omega\] after this one

OpenStudy (anonymous):

@Herp_Derp

OpenStudy (anonymous):

dv/dt=4=angular acceleration v=4t+c (c=0) v=4t=angular velocity tangential velocity=rw =r4t Centripetal acceleration=v^2/r (r4t)^2/r=4 16r t^2=tangential acceleration=r(4) t=sqrt(0.25)

OpenStudy (anonymous):

\[\omega=\alpha t\]

OpenStudy (anonymous):

Ruchi, c was just a possible initial condition, as there is no law that states that particles have to start at rest, so I was just generalising.

OpenStudy (anonymous):

okay bt i hav'nt understood that rw=r4t

OpenStudy (anonymous):

w=angularacceleration x t angular acceleration here=4

OpenStudy (anonymous):

r4t?

OpenStudy (anonymous):

Sorry, tangential velocity=rw= r(4t)